Status Report

NASA Space Station On-Orbit Status 18 Feb 2004

By SpaceRef Editor
February 18, 2004
Filed under , , ,
NASA Space Station On-Orbit Status 18 Feb 2004

All ISS systems continue to function nominally, except as noted previously or below.

Onboard sleep cycle continues with 9:00pm wakeup and 12:30pm bedtime through Friday (2/20).

Before breakfast and physical exercise, CDR/SO Michael Foale and FE Alexander Kaleri conducted the Russian crew health-monitoring program’s medical assessment MO-9/Urinalysis. After the sessions, Kaleri stowed the Urolux equipment. [MO-9 is biochemical urinalysis, conducted regularly every 30 days and also before and after EVAs. It is one of five nominal Russian medical tests adopted by NASA for US crewmembers for PHS evaluation exams (with or without blood labs). The analysis uses the sophisticated in-vitro diagnostic apparatus “Urolux” developed for the Mir program. The device is first calibrated with prepared calibration strips (if not used for more than seven days), then receives the measuring strips with the subject’s urine samples for automatic (photometric) analysis. LEDs indicate immediately if the data are within (green) or outside (red) the physiological norm, and they are also printed on a tape for report to MCC-M/TsUP (actually to IBMP, the Moscow Institute of Biomedical Problems). ]

After breakfast, the crew also completed the MO-5 MedOps protocol of cardiovascular evaluation during graded exercises on the VELO cycle ergometer, each assisting the other in turn. [The assessment uses the Gamma-1 ECG equipment with biomed harness, skin electrodes and a blood pressure and rheoplethysmograph cuff wired to the cycle ergometer’s instrumentation panels. For the graded exercise, Alex and Mike worked the pedals after a prescribed program at load settings of 125, 150, and 175 watts for three minutes each. All measurements were recorded and telemetered during Daily Orbits 2 and 3 to TsUP, where a specialist controlled the workout.]

At 12:40, the crew conducted a teleconference with MCC-Houston via S-band, during which critical details of the upcoming EVA-9 were discussed. [Specific tag-up topics were the new flight rules (FRs) written jointly by U.S. and Russian specialists for this EVA, the use of special EVA gear such as retractable equipment tethers (RETs), the swing arm, wire ties, and the disposable in-suit drink bags (DIDBs), as well as contingency ingress in the U.S. Airlock, which will be configured as a backup ingress path in the unlikely event of multiple failures. The new FRs address EVA termination criteria for specific ISS systems failures (including loss of one of the two U.S. internal thermal cooling loops, a fire on board, depressurization of the ISS, etc.), safety procedures in case of Service Module (SM) thruster firings, should they become necessary, avoidance and clean-up of toxic residues from jet firings, etc.]

Foale and Kaleri reviewed the timeline for tomorrow’s exercise in translating from the DC-1 to the Soyuz orbital module (BO) in Orlan suits, tagging up with ground specialists for discussion. The crew also had more time scheduled for equipment preparations in the DC-1, such as configuring the BK-3 onboard oxygen tank and BNP portable O2 repress bottle, performing height adjustments on the suits, leak and valve functionality checks, etc. Kaleri also had a separate tag-up with specialists on the necessary preparations of the Soyuz BO. [The demo will start tonight at ~11:15pm EST with teardown & removal of the air ducts between the DC-1 and BO, followed by ingress in the Orlan-Ms, checkout of comm & biomedical telemetry via the BSS interface system for vital signs and equipment monitoring, final tests of Soyuz TMA-3 systems, valve functionality, etc. Sealing of the Orlan backpack “doors” is scheduled for about 4:00am tomorrow morning. After more leak checks and a final height fit check at reduced pressure, the actual translation will begin at ~4:40am. Ingress into the Soyuz BO, hatch closing and Orlan doffing inside the BO are estimated to take about an hour. Translation through the narrow DC1-BO hatch opening will be done head first, with one arm extended forward and the other pressed alongside the body. Kaleri will be first, with Foale assisting and pushing if necessary. Sasha then assists Mike, if required, by pulling his hand or putting a pry bar in his gloved hand for squeezing through. The hatch cover will then be closed (which could cause loss of comm with MCC and between the two). FE and CDR then exit from the Orlans one at a time, at ~5:20am. Post-training activities back in the ISS will follow at ~5:35am.]

Sasha attended to his regular daily inspection of the BIO-5 Rasteniya-2 (“Plants-2”) experiment which studies growth and development of plants (peas) under spaceflight conditions in the Lada-4 greenhouse. [Regular daily maintenance of the experiment involves monitoring of seedling growth, humidity measurements, moistening of the substrate if necessary, and photo/video recording.]

Mike Foale transferred exercise data files from the TVIS treadmill to the medical equipment computer (MEC) for subsequent downlink and completed the weekly TVIS maintenance, including the regular inspection of the wire ropes for signs of fraying. He also did the regular (every other week) inspection of the RED (resistive exercise device).

The CDR completed his 13th weekly filling-out of the Food Frequency Questionnaire (FFQ), which keeps a log of his nutritional intake over time on the medical equipment computer (MEC). 

Mike conducted the daily routine maintenance on the SOZh life support system, comprising the water supply equipment, food supply subsystem (SOP), and sanitary hygiene equipment (SGO). He also completed the regular routine status checkup of autonomous Increment 8 payloads.

The crew worked out according to their regular daily physical exercise program on TVIS and CEVIS cycle (aerobic), VELO ergometer with force loader and RED (anaerobic).

At dinnertime (10:30am), as every day, the crew supported the Renal Stone prevention experiment by taking the test medication (either potassium citrate or placebo tablets) until the next sample collection phase in early April this year.

Moscow’s logic change to the Elektron buffer tank to increase supply line pressure (see yesterday’s status report) did not have the desired effect to keep the MNO and MNR micropumps from shutting down. Kaleri took some voltage measurements on the pumps today, but the Elektron O2 generator is still down. 

As planned, several hours ago the cabin atmosphere was repressurized with O2 from Progress 13P, raising ppO2 by 8 mmHg. As shown by the MCA (major constituents analyzer), ppO2 is currently at 162 mmHg.

The Russian SKV-1 air conditioner is operating with the #2 internal heating loop (KOB-2), which collects more condensate than KOB-1, keeping ppH2O (humidity) levels nominally between 7 and 8.5 mmHg.

The crew’s sleep time began at 12:30pm EST, to extend until tonight, 9:00pm, with a very busy day ahead.

Moscow has offered to accept some useful U.S. return cargo as ballast in the Soyuz TMA-3 descent module for a contingency return, should it become necessary.

Today’s CEO targets, in the current LVLH attitude no longer limited by flight rule constraints on the use of the Lab nadir/science window, except for the shutter closure and condensation-prevention plan (limited to 90 min. in 24 hours), were Taiwan Smog (pass across the northern tip of Taiwan:  smog reported in cities only in the southern half of the island. Trying to shoot obliques that capture a margin of the smog mass), Internal waves, Vietnam (looking right toward the glint point for internal waves off the coast of South Vietnam. The zone of mapping interest stretches from Hainan Island [southern China] to Borneo. The glint point crosses the southern half of this zone), Gulf of St Lawrence (Dynamic event. Two opportunities. Looking for cloud streamers as the cold dry air mass flows out over the water and picks up moisture. Cloud is channeled east by the gulf as air flows between the higher coasts. Also looking for newly formed ice), Dust, Niger and Chad (continuing the crew’s documentation of the present major dust event, looking left towards the Tibesti massif from this descending track over Lake Chad. Recent ISS/CEO images show that a wide roughly linear band of lighter surface materials [muds of the ancient floor of paleo-lake Megachad] is the main supplier of dust to the atmosphere during this event. This contradicts accepted geological wisdom that dry lake floors do not supply dust because fine material is cemented by clay and salts. Dust source points [“hotspots” in the jargon] are a major topic of interest in aerosol studies), and Bamako, Mali (this hard-to-see town lies on the Niger River at the head of the Niger inland delta. Transportation routes focus on the town that may also lie in the center of a de-vegetated patch).

CEO images can be viewed at the websites.

See also the website “Space Station Challenge” at

U.S. and Russian Segment Status (as of 2:46pm EST).

Environmental Control and Life Support (ECLSS) and Thermal Control (TCS):

  • Elektron O2 generator is powered Off. Vozdukh CO2 scrubber is On (in Automatic Mode). U.S. CDRA CO2 scrubber is on Standby (ready in dual-bed mode). TCCS (trace contaminant control subsystem) is operating. SM Gas Analyzer has been calibrated and is used for ppO2 and ppCO2 monitoring. MCA (major constituents analyzer) is On to verify CSA-CP readings and support O2 repress (was in Life Extending Mode). BMP Harmful Impurities unit: absorbent bed #1 in Purify mode, bed #2 in Purify mode. RS air conditioner SKV-1 is On, SKV-2 is Off (repair now completed; to be tested ASAP).
  • SM Working Compartment:  Pressure (mmHg) — 742; temperature (deg C) — 26.0; ppO2 (mmHg) — 146.8; ppCO2 (mmHg) — 2.5;
  • SM Transfer Compartment:  Pressure (mmHg) — 751; temperature (deg C) — 20.7.
  • FGB Cabin:  Pressure (mmHg) — 744; temperature (deg C) — 22.7.
  • Node:  Pressure (mmHg) — 742.62; temperature (deg C) — 23.5 (shell); ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • U.S. Lab:  Pressure (mmHg) — 744.72; temperature (deg C) — 24.8; ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • Joint Airlock (Equip. Lock):   Pressure (mmHg) — 744.82; temperature (deg C) — 23.5; shell heater temp (deg C) — 24.2, ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • PMA-1:  Shell heater temp (deg C) — 22.8
  • PMA-2:  Shell heater temp (deg C) — 11.3

(n/a = data not available)

Electrical Power Systems (EPS):

  • Both P6 channels fully operational. BGA (beta gimbal assembly) 2B and 4B both in Directed Position (2B: 235 deg; 4B: 125 deg); non-suntracking, “night glider”/”sun slicer” drag reduction mode.
  • SM batteries:  Battery #8, formerly known as #7, is failed (to be replaced); all other batteries (7) are in “Partial Charge” mode.  
  • FGB batteries:  Battery #3 is off (capacity restoration mode, ROM); battery #5 is in “Cycle” mode; all other batteries (4) are in “Partial Charge” mode.
  • Plasma Contactor Unit PCU-1 in Standby mode; PCU-2 in Standby mode.

Command & Data Handling Systems:

  • C&C-1 MDM is prime, C&C-2 is back-up, and C&C-3 is in standby.
  • GNC-1 MDM is prime; GNC-2 is Backup.
  • INT-2 is operating; INT-1 is Off.
  • EXT-2 is On (primary), EXT-1 is Off (both now upgraded to R3).
  • LA-1, LA-2 and LA-3 MDMs are all operating.
  • PL-1 MDM is Off; PL-2 MDM is Operational.
  • APS-1 (automated payload switch #1) and APS-2 are both On.
  • SM Terminal Computer (TVM): 2 redundant lanes (of 3) operational (string 1 dropped out 11/22).
  • SM Central Computer (TsVM): 2 redundant lanes (of 3) operational (string #3 dropped out 10/22).

Propulsion System:

  • Total propellant load available: 4070 kg (8972 lb) as of 2/12/04  [SM(755) + FGB(2656) + Progress M(0) + Progress M-1(659)]. (Capability: SM — 860 kg; FGB — 6120 kg).

Attitude Control Systems:

  • 3 CMGs on-line (CMG-1 failed).
  • State vector source — U.S. SIGI-1 (GPS)
  • Attitude source — U.S. SIGI-1 (GPS)
  • Angular rate source — RGA-1

Flight Attitude:

  • LVLH YVV (local vertical/local horizontal = “earth-fixed”: z-axis in local vertical, y-axis in velocity vector [yaw: -90 deg, pitch: -9 deg, roll: 1.7 deg]), with CMG/TA (thruster assist) Momentum Management.

Communications & Tracking Systems:

  • FGB MDM-1 is powered Off; FGB MDM-2 is operational.
  • All other Russian communications & tracking systems are nominal.
  • *S-band is operating nominally (on string 2).
  • Ku-band is operating nominally.
  • Audio subsystem is operating nominally (IAC-1 is prime, IAC-2 is off).
  • Video subsystem operating nominally.
  • HCOR (high-rate communications outage recorder) is operating nominally.


  • SSRMS/Canadarm2 based at Lab PDGF/LEE A, with Keep Alive (KA) power on both strings.
  • MBS: KA power on both strings. 
  • MT: latched and mated at WS4. 
  • POA: KA power on both strings.
  • RWS (robotics workstations): Lab RWS is On (DCP connected); Cupola RWS is Off.

ISS Orbit  (as of this morning, 7:46am EST [= epoch]):

  • Mean altitude — 366.8 km
  • Apogee — 371.4km
  • Perigee — 362.1 km
  • Period — 91.88 min.
  • Inclination (to Equator) —  51.628 deg
  • Eccentricity — 0.0006898
  • Orbits per 24-hr. day — 15.67
  • Mean altitude loss last 24 hours — 80 m
  • Revolutions since FGB/Zarya launch (Nov. ’98)  — 29954

For more on ISS orbit and worldwide ISS naked-eye visibility dates/times, see

SpaceRef staff editor.